A concrete mix is in volume proportions of 1 part cement, 2 parts water, 2 parts aggregate, and 3 parts sand. How many cubic feet of each ingredient are needed to make 54 cu ft of concrete?
Answer by edjones(4080)

1 part cement, 2 parts water, 2 parts aggregate, and 3 parts sand
1+2+2+3=8
54/8=6.75
.
6.75*2=13.5
6.75*3=20.25
.
6.75 ft^3 cement
13.5 ft^3 water
13.5 ft^3 aggregate
20.25 ft^3 sand
------------------add
54 ft^3 concrete
.
Ed